# HiM2SAM: Enhancing SAM2 with Hierarchical Motion Estimation and Memory Optimization towards Long-term Tracking

### v2

@@ Abstract
Visual object tracking (VOT) in open-world environments remains a fundamental yet challenging problem, particularly in scenarios involving complex motion, target disappearance and reappearance, and long-term tracking. Traditional tracking-by-detection methods (e.g. [26][27][33]) often rely on predefined object classes and motion models such as linear Kalman Filters [12], limiting their adaptability to the diverse and complex dynamics of real-world scenarios.
@@
HiM2SAM remains training-free, incurs minimal additional cost, and delivers performance gains across VOT benchmarks, particularly on LaSOT [7][8] series datasets. It also significantly boosts performance on smaller SAM2 variants, highlighting its effectiveness in efficient, long-term object tracking. An overview of our method and its comparison with SOTA performance on LaSOT are shown in Figure 1.
@@ 2 Related Work
Segment Anything Model 2 (SAM2). SAM2 [22] extends the Segment Anything Model (SAM) [16] to video by introducing a memory bank that stores previous frame features and masks. During inference, cross-attention is computed between the current frame and memory frames to produce segmentation masks without requiring prompts. Despite strong performance, SAM2’s limited memory capacity and lack of temporal fusion hinder its robustness against distractors and long-term occlusions. To mitigate this, SAMURAI [32] adds motion prediction and mask reweighting using Kalman filters; SAM2Long [5] treats mask selection as a sequential decision-making task; and DAM4SAM [30] identifies distractor frames and leverages them to guide long-term memory updates, retaining high-confidence features for improved tracking stability.
Point Tracking. Tracking points across video frames supports fine-grained tasks like 3D reconstruction and motion analysis, focusing on pixel-level correspondence rather than object semantics. Recent works such as PIPs [9], TAPIR [6], and CoTracker [15] introduce differentiable correlation, global matching, and attention-based mechanisms to improve robustness. CoTracker3 [14] simplifies this paradigm by combining efficient 4D correlation and iterative track updates, achieving SOTA accuracy with significantly less training data. Its ability to model dense, long-term motion complements segmentation-based approaches, making it especially useful in complex, dynamic scenes.
@@ 3 Revisit Segment Anything Model 2
SAM2 builds upon the architecture of SAM (Segment Anything Model) and extends it to video object segmentation by introducing a memory mechanism, which enables prompt-free segmentation across frames. In addition to a Vision Transformer-based image encoder [23] that extracts per-frame visual features, the memory mechanism introduces several key components tailored for video segmentation: a mask decoder, a memory encoder and memory bank, and memory attention. The structure of SAM2 is shown in upper part in Fig.2
@@ 4 Proposed Method
4.1 Hierarchy Motion Estimation
@@ 5.1 Benchmarks
| LaSOT | LaSOT | |||||
| Tracker | AUC(%) | P(%) | P(%) | AUC(%) | P(%) | P(%) |
@@ 5.2 Quantitative Results
Overal Results.
We compare state-of-the-art supervised tracking methods and zero-shot trackers. Unless otherwise specified, all SAM2-based experiments in this paper refer to the latest SAM2.1 model. To ensure fair comparison, only methods capable of achieving real-time performance are considered. Table LABEL:tab:main_results presents the quantitative results on LaSOT series. Our proposed method achieves competitive overall performance among zero-shot methods. Despite the absence of training data, HiM2SAM demonstrates significant gains over supervised approaches on LaSOT, showcasing the strong generalization ability.
Furthermore, our method consistently outperforms baseline approaches across all SAM2 model sizes. The gains are especially notable in smaller models, which often struggle with complex tracking due to limited capacity. This demonstrates the value of our hierarchical motion estimation and long-short memory mechanism. By enhancing frame correspondence and managing memory efficiently, these components improve tracking robustness with minimal computational overhead, making them particularly beneficial for lightweight models.
As shown in Table LABEL:tab:vot, we evaluate zero-shot performance on other two long-term tracking benchmarks: VOT-LT2020 and VOT-LT2022 using the large-size models. Compared with previous zero-shot methods, HiM2SAM achieves the highest F-score on both datasets, indicating stronger robustness under long-term tracking scenarios. These results suggest that our design, which incorporates hierarchical motion estimation and memory management, is effective in handling frequent target disappearances and complex scene variations.
Tracking Attributes.
We further conduct an attribute-based analysis on both LaSOT and LaSOT using the base-size model, focusing on AUC performance. The attributes defined in [8] can be categorized into two categories, Dynamic Motion and Degraded Visibility. As shown in Table 3, our method consistently outperforms SAMURAI across all attributes, indicating strong generalization to diverse and challenging tracking scenarios.
@@ 5.3 Ablation Studies and Runtime Analysis
Results in Table LABEL:tab:abl show that each module improves tracking over the SAM2 baseline, with the full combination yielding the best performance and only modest runtime overhead. Runtime is measured on a single NVIDIA A6000 GPU and Intel i7-9700K CPU. Latency is averaged over all frames from a representative LaSOT subset. Notably, using only PT achieves higher accuracy than KF but at a significant computational cost. In contrast, HiM2SAM’s hierarchical motion estimation combines KF’s efficiency and PT’s precision, minimizing runtime overhead. These results confirm the modules’ complementarity and validate HiM2SAM’s design as a fast and accurate long-term tracker.
@@ 5.4 Sensitivity to Parameters
We conduct a parameter sensitivity analysis on the LaSOT dataset to examine the robustness of our tracker HiM2SAM-L under different settings. Figure 3 presents the evaluation results for three key parameters: (1) the number of frames used by the point tracker for temporal propagation; (2) the IoU threshold used during coarse motion estimation to decide whether a fine motion estimation is needed (triggered when the IoU falls below ); and (3) different combinations of frame intervals for long-term memory (LM) and short-term memory (SM). We find that while the tracker maintains overall stability, careful tuning of and memory intervals can yield marginal improvements in AUC.
@@ 5.5 Qualitative Results
We compare HiM2SAM with SAMURAI and DAM4SAM in challenging long-term tracking scenarios, as shown in Fig. 4. These sequences involve cluttered backgrounds, occlusions, and significant appearance variations. Our pixel-level motion estimation enables more precise mask boundaries, particularly under shape deformation. Meanwhile, the long-short memory mechanism enhances temporal consistency by improving robustness to occlusion and appearance changes. Together, these modules contribute to more stable and continuous tracking over long periods.

### v3

@@ Abstract
Visual object tracking (VOT) in open-world environments remains a fundamental yet challenging problem, particularly in scenarios involving complex motion, target disappearance and reappearance, and long-term tracking. Traditional tracking-by-detection methods (e.g., [26, 27, 33]) often rely on predefined object classes and motion models such as linear Kalman filters [11], limiting their adaptability to the diverse and complex dynamics of real-world scenarios.
@@
HiM2SAM remains training-free, incurs minimal additional cost, and delivers performance gains across VOT benchmarks, particularly on LaSOT [7, 8] series datasets. It also significantly boosts performance on smaller SAM2 variants, highlighting its effectiveness in efficient, long-term object tracking. An overview of our method and its comparison with SOTA performance on LaSOT are shown in Figure 1.
@@ 2 Related Work
Segment Anything Model 2 (SAM2). SAM2 [22] extends the Segment Anything Model (SAM) [15] to video by introducing a memory bank that stores previous frame features and masks. During inference, cross-attention is computed between the current frame and memory frames to produce segmentation masks without requiring prompts. Its capabilities have been validated in various tasks [36, 5, 18]. Despite strong performance, SAM2’s limited memory capacity and lack of temporal fusion hinder its robustness against distractors and long-term occlusions. To mitigate this, SAMURAI [32] adds motion prediction and mask reweighting using Kalman filters; SAM2Long [6] treats mask selection as a sequential decision-making task; and DAM4SAM [30] identifies distractor frames and leverages them to guide long-term memory updates, retaining high-confidence features for improved tracking stability.
Point Tracking. Tracking points across video frames supports fine-grained tasks like 3D reconstruction and motion analysis, focusing on pixel-level correspondence rather than object semantics. Recent works such as CoTracker [14] introduce differentiable correlation, global matching, and attention-based mechanisms to improve robustness. CoTracker3 [13] simplifies this paradigm by combining efficient 4D correlation and iterative track updates, achieving SOTA accuracy with significantly less training data. Its ability to model dense, long-term motion complements segmentation-based approaches, making it especially useful in complex, dynamic scenes.
@@ 3 Revisit Segment Anything Model 2
SAM2 builds upon the architecture of SAM (Segment Anything Model) and extends it to video object segmentation by introducing a memory mechanism, which enables prompt-free segmentation across frames. In addition to a Vision Transformer-based image encoder [23] that extracts per-frame visual features, the memory mechanism introduces several key components tailored for video segmentation: a mask decoder, a memory encoder and memory bank, and memory attention. The structure of SAM2 is shown in upper part in Figure 2.
@@ 4 Proposed Method
4.1 Hierarchical Motion Estimation
@@ 5.1 Benchmarks
| Trackers | LaSOT | LaSOT | ||||
| AUC(%) | P(%) | P(%) | AUC(%) | P(%) | P(%) | |
@@ 5.2 Quantitative Results
Overal Results. We compare state-of-the-art supervised tracking methods and zero-shot trackers. Unless otherwise specified, all SAM2-based experiments in this paper refer to the latest SAM2.1 model. For our method, CoTracker3 [13] is employed for point propagation. To ensure fair comparison, only methods capable of achieving real-time performance are considered. Table 1 presents the quantitative results on LaSOT series. Our proposed method achieves competitive overall performance among zero-shot methods. Despite the absence of training data, HiM2SAM demonstrates significant gains over supervised approaches on LaSOT, showcasing the strong generalization ability.
Furthermore, our method consistently outperforms baseline approaches across all SAM2 model sizes. The gains are especially notable in smaller models, which often struggle with complex tracking due to limited capacity. This demonstrates the value of our hierarchical motion estimation and long-short memory mechanism. By enhancing frame correspondence and managing memory efficiently, these components improve tracking robustness with minimal computational overhead, making them particularly beneficial for lightweight models. Importantly, these improvements are achieved without sacrificing the class-agnostic nature of SAM2, demonstrating that our approach maintains strong robustness and generalization across diverse objects and scenarios.
As shown in Table 2, we evaluate zero-shot performance on other two long-term tracking benchmarks: VOT-LT2020 and VOT-LT2022 using the large-size models. Compared with previous zero-shot methods, HiM2SAM achieves the highest F-score on both datasets, indicating stronger robustness under long-term tracking scenarios. These results suggest that our design, which incorporates hierarchical motion estimation and memory management, is effective in handling frequent target disappearances and complex scene variations.
Tracking Attributes. We further conduct an attribute-based analysis on both LaSOT and LaSOT using the base-size model, focusing on AUC performance. The attributes defined in [8] can be categorized into two categories, Dynamic Motion and Degraded Visibility. As shown in Table 3, our method consistently outperforms SAMURAI across all attributes, indicating strong generalization to diverse and challenging tracking scenarios.
@@ 5.3 Ablation Studies and Runtime Analysis
Results in Table LABEL:tab:abl show that each module improves tracking over the SAM2 baseline, with the full combination yielding the best performance and only modest runtime overhead. Runtime is measured on a single NVIDIA A6000 GPU and Intel i7-9700K CPU. Latency is averaged over all frames from a representative LaSOT subset. Notably, using only PT achieves higher accuracy than KF but at a significant computational cost. In contrast, HiM2SAM’s hierarchical motion estimation combines KF’s efficiency and PT’s precision, minimizing runtime overhead. The last row reports our final result, which demonstrates the complementarity of the modules and preserves the real-time capability of the original SAM2 model, validating HiM2SAM’s design as a fast and accurate long-term tracker.
@@ 5.4 Sensitivity to Parameters
We conduct a parameter sensitivity analysis on the LaSOT dataset to examine the robustness of our tracker HiM2SAM-L under different settings. Figure 3 presents the evaluation results for three key parameters: (1) the number of frames used by the point tracker for temporal propagation; (2) the IoU threshold used during coarse motion estimation to decide whether a fine motion estimation is needed (triggered when the IoU falls below ); and (3) different combinations of frame intervals for long-term memory (LM) and short-term memory (SM). We find that the tracker maintains overall stability across a wide range of parameter settings. Importantly, HiM2SAM consistently outperforms the SAM2 baseline, demonstrating the robustness and effectiveness of our method despite the presence of tunable parameters. The final selection of parameters is provided in our released code.
@@ 5.5 Qualitative Results
We compare HiM2SAM with SAMURAI and DAM4SAM in challenging long-term tracking scenarios, as shown in Figure 4. These sequences involve cluttered backgrounds, occlusions, and significant appearance variations. Our pixel-level motion estimation enables more precise mask boundaries, particularly under shape deformation. Meanwhile, the long-short memory mechanism enhances temporal consistency by improving robustness to occlusion and appearance changes. Together, these modules contribute to more stable and continuous tracking over long periods.